# SynLlama: Generating Synthesizable Molecules and Their Analogs with Large Language Models

### v4

@@ 3.2 Synthesizable Analog Search for De Novo Molecules
Previous research has shown that molecules proposed by generative models for binding to protein targets often face challenges in both reliability and practical synthesizability28, 29, 30. In particular, medicinal chemists are often reluctant to devote time and expensive resources to specialized synthesis of hit molecules due to the high false positive rates arising from the unreliability of docking scores in drug discovery. Instead finding closely related compounds that are constrained to Enamine BBs allows for an inexpensive purchase to verify hits before further refinements are deployed to gain lead drug molecules. In this section, we demonstrate SynLlama’s potential to bridge the gap between generative molecule design and practical synthesis for molecules that are optimized for drug-like applications such as binding assays.
@@
Specifically, we consider two different generative methods for de novo drug molecules in Figure 2: a 1D-to-3D LSTM model, iMiner19, which optimizes drug-likeness and AutoDock-vina73 docking scores, and Pocket2Mol22, a 3D graph transformer model which also optimizes AutoDock-vina73 docking scores. Using iMiner we generated 500 molecules for the SARS-CoV-2 Main Protease (SARS2 MPro)74, and used Pocket2Mol to generate 500 molecules each for the protein targets Thrombin75 and TYK276, 77. The first interesting observation is that only 1% of generated molecules from both iMiner and Pocket2Mol can be synthetically reconstructed using SynLlama or ChemProjector, emphasizing that the generative models create difficult synthesis targets. Hence all 500 compounds for each protein target were then processed through SynLlama trained on RXN 2 to generate synthesizable analogs constrained to Enamine BBs. For all generated analogs we performed molecular docking with AutoDock-vina via using the same protocol as the original binders.
@@
Figure 2(a,b) shows the RMSE of the docking scores between the target compounds and the generated analogs are 1.44 kcal/mol and 1.11 kcal/mol for Pocket2Mol and iMiner, respectively, both of which are within the acceptable range of inherent docking score errors reported by Trott et al73. Furthermore, as demonstrated in Figures 2(c,d), the SA score distribution of the SynLlama analogs generated for thrombin and TYK2 from Pocket2Mol show a notable improvement in synthetic accessibility at the expense of reduced similarity below the benchmarks in Table 2. SynLlama slightly improves SA for iMiner generated compounds for SARS2 MPro while maintaining a good similarity score on par with the benchmarks in Table 2. The fact that iMiner uses a drug-likeness score as part of its loss function19 may explain its better performance, or perhaps SARS2 MPro is an easier protein target than Pocket2Mol’s thrombin and TYK2 protein targets.
@@
Figure 2: SynLlama performance on generating synthesizable analogs for Pocket2Mol and iMiner proposed binders of SARS2 MPro74, Thrombin75, and TYK276, 77. Correlation plot comparing docking scores of (a) Pocket2Mol and (b) iMiner generated molecules and the average Vina docking scores of ten most similar analogs from SynLlama trained with RXN 2. Each data point is color-coded by the average Morgan fingerprint similarity computed between the generated and analog molecules. The shaded area represents an energy uncertainty range of $\pm 2kcal/mol$ for docking73. (c) Synthetic accessibility (SA) score distribution of Pocket2Mol, iMiner, and unsynthesizable molecules and SynLlama-proposed analogs. iMiner analogs generated with SynLlama trained on RXN 1 showed similar results as reported in Supplementary Figure S6. The kernel density in Supplementary Figure S7 further confirms our finding that the analogs consistently shift toward better SA without undermining the overall docking score distribution. (d) average Morgan fingerprint similarity score between the target molecules and their top-10 proposed analogs.
@@
The final comparison consists of molecules that were identified as unsynthesizable by Gao et al30 via ASKCOS48. As also seen in Figure 2, the similarity scores of SynLlama analogs for this set are better than that observed for the Pocket2Mol generative model, and the generated analogs show a significant decrease in SA score, representing SynLlamas effective strategy of improving synthetic accessibility via analog generation. In Supplementary Figure S8, we also highlight the few concrete examples of target-analog pairs where objective scores are maintained while target synthetic accessibility scores are substantially reduced. Overall, these collective results highlight SynLlama’s utility in effectively proposing synthesizable analogs for de novo molecules, thereby enhancing their synthetic accessibility without compromising their desired drug-related properties.
@@ 3.3 Local Hit Expansion for Binder Molecules
scaffold rather than only on a whole target molecule. In a final task, we apply SynLlama to expand on hit molecules for three protein targets used in the previous section, SARS2 Mpro78, Thrombin75, and TYK276, 77 to discover synthesizable molecules that have better relative binding free energies (RBFEs) confirmed by both experiments and accurate free energy perturbation (FEP) calculations.
As shown in Figure 4(a), the hit molecule for SARS2 Mpro (7LTJ_Lead) has a core scaffold of uracil and ortho-dichlorobenzene connected by a piperizine linker. Inspired by an experimental hit expansion campaign by Kneller et al. 79, we follow their practice to propose only functional group substitutions on the benzene ring while keeping the linker and uracil intact. In Figures 4(c) and 4(e), we use the best-performing molecules from the Schrodinger FEP benchmarking set80 as our hits for Thrombin and TYK2. To expand on these hit molecules, we first identify the maximum common scaffolds among each group of molecules in the FEP benchmarking set and use the identified scaffolds to guide our selection of analog molecules. Specifically, we use SynLlama to generate 50 synthesizable analogs constrained to only Enamine BBs of the hit compound and filter for molecules that retain the scaffold. In the end, we harvest a total of 8, 14, and 11 analog molecules that fulfill the criteria for SARS2 Mpro, Thrombin, and TYK2, respectively. The analogs are then placed in a pose configuration similar to the original hit molecule for downstream FEP calculations.
@@
Moreover, the fact that all suggested analogs also come with predicted pathways using common reaction templates and purchasable BBs from Enamine suggests SynLlama’s practical use for drug discovery. Because of this composite capacity of optimizing both potency and synthetic accessibility for small molecules, SynLlama successfully rediscovered Analog_1, the most potent compound reported by Kneller et al. in their hit expansion campaign for SARS2 Mpro79. Furthermore, for the Thrombin case, SynLlama explores the R2 substitution site, a region previously unaddressed by earlier molecular series, and demonstrates the generation of more potent molecules via FEP. These results confirm that SynLlama effectively explores local chemistry with readily available building blocks, providing a direct and efficient path for medicinal chemists to accelerate hit expansion.
@@ 4 Discussion and Conclusions
Motivated by recent advances in LLMs for chemistry60, 58, 27, 59, we aim to leverage data-efficient supervised fine-tuning (SFT) to transform the general-purpose Meta Llama 3 into SynLlama, an LLM-based generator capable of proposing synthesizable molecules and deducing synthetic routes for target molecules or their close analogs. Throughout the study, we successfully show that SynLlama can effectively explore a custom-defined chemical search space composed of around 230,000 Enamine building blocks (BBs) and well-validated organic reactions (RXNs), after it has been fine-tuned on synthetic pathway data sampled from this specified chemical space. What’s more, despite utilizing nearly two orders of magnitude fewer synthetic pathways in training, SynLlama exhibits strong performance in key drug discovery tasks compared to existing models. Specifically, we have demonstrated that SynLlama can effectively aid in various stages of drug discovery that include synthesis planning, synthesizable analog generation for de novo molecules, and local hit expansion.
Because SynLlama is built on a general-purpose LLM instead of training from scratch54, 55, 56, it offers a number of unique advantages and possibilities for further improvement. For example, when generating our fine-tuning data, we sampled from the predefined chemical search space of 230K Enamine building blocks (BBs) and two sets of reaction templates (RXNs), but we did not embed these extensive requirements in the context window of the LLM. As a result, for our largest dataset with only 2 million synthetic pathways, the model only saw each BB a few dozen times, while each RXN template appeared hundreds of thousands of times. Consequently, while SynLlama efficiently memorizes the allowed RXNs, it only captures the distribution of Enamine BBs, which enables SynLlama to extrapolate to unseen yet purchasable building blocks outside of Enamine. This generative ability surpasses other existing methods such as ChemProjector and Synformer that can only explore a predefined building block search space, such as within the Enamine Diversity Set9, and limits their ability to propose alternative synthesis pathway with novel building blocks.

### v5

@@ 3.2 Synthesizable Analog Search for De Novo Molecules
Previous research has shown that molecules proposed by generative models for binding to protein targets often face challenges in both reliability and practical synthesizabilitysumita_hunting_2018, zhavoronkov_deep_2019, gao_synthesizability_2020. In particular, medicinal chemists are often reluctant to devote time and expensive resources to specialized synthesis of hit molecules due to the high false positive rates arising from the unreliability of docking scores in drug discovery. Instead finding closely related compounds that are constrained to Enamine BBs allows for an inexpensive purchase to verify hits before further refinements are deployed to gain lead drug molecules. In this section, we demonstrate SynLlama’s potential to bridge the gap between generative molecule design and practical synthesis for molecules that are optimized for drug-like applications such as binding assays.
@@
Specifically, we consider two different generative methods for de novo drug molecules in Figure 2: a 1D-to-3D LSTM model, iMineriminer, which optimizes drug-likeness and AutoDock-vinaTrott2010 docking scores, and Pocket2Moldenovo_pocket2mol, a 3D graph transformer model which also optimizes AutoDock-vinaTrott2010 docking scores. Using iMiner we generated 500 molecules for the SARS-CoV-2 Main Protease (SARS2 MPro)zhang2021potent, and used Pocket2Mol to generate 500 molecules each for the protein targets Thrombinthrombin and TYK2tyk2, tyk2_2. The first interesting observation is that only 1% of generated molecules from both iMiner and Pocket2Mol can be synthetically reconstructed using SynLlama or ChemProjector, emphasizing that the generative models create difficult synthesis targets. Hence all 500 compounds for each protein target were then processed through SynLlama trained on RXN 2 to generate synthesizable analogs constrained to Enamine BBs. For all generated analogs we performed molecular docking with AutoDock-vina via using the same protocol as the original binders.
@@
Figure 2(a,b) shows the RMSE of the docking scores between the target compounds and the generated analogs are 1.44 kcal/mol and 1.11 kcal/mol for Pocket2Mol and iMiner, respectively, both of which are within the acceptable range of inherent docking score errors reported by Trott et alTrott2010. Furthermore, as demonstrated in Figures 2(c,d), the SA score distribution of the SynLlama analogs generated for thrombin and TYK2 from Pocket2Mol show a notable improvement in synthetic accessibility at the expense of reduced similarity below the benchmarks in Table 2. SynLlama slightly improves SA for iMiner generated compounds for SARS2 MPro while maintaining a good similarity score on par with the benchmarks in Table 2. The fact that iMiner uses a drug-likeness score as part of its loss functioniminer may explain its better performance, or perhaps SARS2 MPro is an easier protein target than Pocket2Mol’s thrombin and TYK2 protein targets.
@@
Figure 2: SynLlama performance on generating synthesizable analogs for Pocket2Mol and iMiner proposed binders of SARS2 MProzhang2021potent, Thrombinthrombin, and TYK2tyk2, tyk2_2. Correlation plot comparing docking scores of (a) Pocket2Mol and (b) iMiner generated molecules and the average Vina docking scores of ten most similar analogs from SynLlama trained with RXN 2. Each data point is color-coded by the average Morgan fingerprint similarity computed between the generated and analog molecules. The shaded area represents an energy uncertainty range of $\pm 2kcal/mol$ for dockingTrott2010. (c) Synthetic accessibility (SA) score distribution of Pocket2Mol, iMiner, and unsynthesizable molecules and SynLlama-proposed analogs. iMiner analogs generated with SynLlama trained on RXN 1 showed similar results as reported in Supplementary Figure S6. The kernel density in Supplementary Figure S7 further confirms our finding that the analogs consistently shift toward better SA without undermining the overall docking score distribution. (d) average Morgan fingerprint similarity score between the target molecules and their top-10 proposed analogs.
@@
The final comparison consists of molecules that were identified as unsynthesizable by Gao et algao_synthesizability_2020 via ASKCOScoley_robotic_2019. As also seen in Figure 2, the similarity scores of SynLlama analogs for this set are better than that observed for the Pocket2Mol generative model, and the generated analogs show a significant decrease in SA score, representing SynLlamas effective strategy of improving synthetic accessibility via analog generation. In Supplementary Figure S8, we also highlight the few concrete examples of target-analog pairs where objective scores are maintained while target synthetic accessibility scores are substantially reduced. Overall, these collective results highlight SynLlama’s utility in effectively proposing synthesizable analogs for de novo molecules, thereby enhancing their synthetic accessibility without compromising their desired drug-related properties.
@@ 3.3 Local Hit Expansion for Binder Molecules
scaffold rather than only on a whole target molecule. In a final task, we apply SynLlama to expand on hit molecules for three protein targets used in the previous section, SARS2 Mpro7ltj, Thrombinthrombin, and TYK2tyk2, tyk2_2 to discover synthesizable molecules that have better relative binding free energies (RBFEs) confirmed by both experiments and accurate free energy perturbation (FEP) calculations.
As shown in Figure 4(a), the hit molecule for SARS2 Mpro (7LTJ_Lead) has a core scaffold of uracil and ortho-dichlorobenzene connected by a piperizine linker. Inspired by an experimental hit expansion campaign by Kneller et al. kneller_hit_expansion_7ltj, we follow their practice to propose only functional group substitutions on the benzene ring while keeping the linker and uracil intact. In Figures 4(c) and 4(e), we use the best-performing molecules from the Schrodinger FEP benchmarking setschrodinger_bnechmark_github as our hits for Thrombin and TYK2. To expand on these hit molecules, we first identify the maximum common scaffolds among each group of molecules in the FEP benchmarking set and use the identified scaffolds to guide our selection of analog molecules. Specifically, we use SynLlama to generate 50 synthesizable analogs constrained to only Enamine BBs of the hit compound and filter for molecules that retain the scaffold. In the end, we harvest a total of 8, 14, and 11 analog molecules that fulfill the criteria for SARS2 Mpro, Thrombin, and TYK2, respectively. The analogs are then placed in a pose configuration similar to the original hit molecule for downstream FEP calculations.
@@
Moreover, the fact that all suggested analogs also come with predicted pathways using common reaction templates and purchasable BBs from Enamine suggests SynLlama’s practical use for drug discovery. Because of this composite capacity of optimizing both potency and synthetic accessibility for small molecules, SynLlama successfully rediscovered Analog_1, the most potent compound reported by Kneller et al. in their hit expansion campaign for SARS2 Mprokneller_hit_expansion_7ltj. Furthermore, for the Thrombin case, SynLlama explores the R2 substitution site, a region previously unaddressed by earlier molecular series, and demonstrates the generation of more potent molecules via FEP. These results confirm that SynLlama effectively explores local chemistry with readily available building blocks, providing a direct and efficient path for medicinal chemists to accelerate hit expansion.
@@ 4 Discussion and Conclusions
Motivated by recent advances in LLMs for chemistryyu2024llasmol, boiko_autonomous_2023, cavanagh_smileyllama_2024, m_bran_augmenting_2024, we aim to leverage data-efficient supervised fine-tuning (SFT) to transform the general-purpose Meta Llama 3 into SynLlama, an LLM-based generator capable of proposing synthesizable molecules and deducing synthetic routes for target molecules or their close analogs. Throughout the study, we successfully show that SynLlama can effectively explore a custom-defined chemical search space composed of around 230,000 Enamine building blocks (BBs) and well-validated organic reactions (RXNs), after it has been fine-tuned on synthetic pathway data sampled from this specified chemical space. What’s more, despite utilizing nearly two orders of magnitude fewer synthetic pathways in training, SynLlama exhibits strong performance in key drug discovery tasks compared to existing models. Specifically, we have demonstrated that SynLlama can effectively aid in various stages of drug discovery that include synthesis planning, synthesizable analog generation for de novo molecules, and local hit expansion.
Because SynLlama is built on a general-purpose LLM instead of training from scratchgao_amortized_2022, luo_projecting_2024, gao_generative_2024, it offers a number of unique advantages and possibilities for further improvement. For example, when generating our fine-tuning data, we sampled from the predefined chemical search space of 230K Enamine building blocks (BBs) and two sets of reaction templates (RXNs), but we did not embed these extensive requirements in the context window of the LLM. As a result, for our largest dataset with only 2 million synthetic pathways, the model only saw each BB a few dozen times, while each RXN template appeared hundreds of thousands of times. Consequently, while SynLlama efficiently memorizes the allowed RXNs, it only captures the distribution of Enamine BBs, which enables SynLlama to extrapolate to unseen yet purchasable building blocks outside of Enamine. This generative ability surpasses other existing methods such as ChemProjector and Synformer that can only explore a predefined building block search space, such as within the Enamine Diversity Setenamine_bb, and limits their ability to propose alternative synthesis pathway with novel building blocks.
@@ 8 SUPPLEMENTARY INFORMATION
TOC Graphic
Synopsis: Fine-tuning on synthetic reactions from commercial building blocks and high-fidelity reactions creates a versatile LLM, SynLlama, for key drug discovery tasks.